# End-to-end Autonomous Driving: Challenges and Frontiers

### v1

@@ Abstract
The autonomous driving community has witnessed a rapid growth in approaches that embrace an end-to-end algorithm framework, utilizing raw sensor input to generate vehicle motion plans, instead of concentrating on individual tasks such as detection and motion prediction. End-to-end systems, in comparison to modular pipelines, benefit from joint feature optimization for perception and planning. This field has flourished due to the availability of large-scale datasets, closed-loop evaluation, and the increasing need for autonomous driving algorithms to perform effectively in challenging scenarios. In this survey, we provide a comprehensive analysis of more than 250 papers, covering the motivation, roadmap, methodology, challenges, and future trends in end-to-end autonomous driving. We delve into several critical challenges, including multi-modality, interpretability, causal confusion, robustness, and world models, amongst others. Additionally, we discuss current advancements in foundation models and visual pre-training, as well as how to incorporate these techniques within the end-to-end driving framework. To facilitate future research, we maintain an active repository that contains up-to-date links to relevant literature and open-source projects at https://github.com/OpenDriveLab/End-to-end-Autonomous-Driving.
@@ 1 Introduction
Conventional autonomous driving systems adopt a modular deployment strategy, wherein each functionality, such as perception, prediction, and planning, is individually developed and integrated into the onboard vehicle. The planning or control module, responsible for generating steering and acceleration outputs, plays a crucial role in determining the driving experience. The most common approach for planning in modular pipelines involves using sophisticated rule-based designs, which are often ineffective in addressing the vast number of situations that occur while driving. Therefore, there is a growing trend to leverage large-scale data and to use learning-based planning as a viable alternative.
@@
In this survey, we conduct an extensive review of this emerging topic. Fig. 1 provides an overview of our work. We begin by discussing the motivation and roadmap for end-to-end autonomous driving systems. End-to-end approaches can be broadly classified into imitation and reinforcement learning, and we provide a brief review of these methodologies. We cover datasets and benchmarks for both closed- and open-loop evaluation. We summarize a series of critical challenges, including interpretability, generalization, world models, causal confusion, and others. We conclude by discussing future trends that we think should be embraced by the community to incorporate the latest developments from data engines, large foundation models, and vehicle-to-everything, amongst others.
Figure 1: Our Survey at a glance. I. Pipeline and methods. We define end-to-end autonomous driving as a learning-based algorithm framework with raw sensor input and planning/control output. We classify the 250+ papers surveyed into imitation learning (IL) and reinforcement learning (RL). II. Benchmarking. We group popular benchmarks into closed-loop and open-loop evaluation, respectively. We cover various aspects of closed-loop simulation and the limitations of open-loop evaluation for this problem. III. Challenges. This is the main section of our work. We list key challenges from a wide range of topics and extensively analyze why these concerns are crucial. We also comment on promising resolutions to these challenges. IV. Future trends. We discuss how end-to-end autonomous driving could benefit from the aid of the rapid development of foundation models, visual pre-training, etc. Some photos by courtesy of online resources.
@@ 1.1 Motivation of an End-to-end System
In the classical pipeline, each model serves a standalone component and corresponds to a specific task (e.g., traffic light detection). Such a design is beneficial in terms of interpretability, verifiability, and ease of debugging. However, since the optimization goal across modules is different, with detection in perception pursuing mean average precision (mAP) while planning aiming for driving safety and comfort, the entire system may not be aligned with a unified target, i.e., the ultimate planning/control task. Errors from each module, as the sequential procedure proceeds, could be compounded and result in an information loss for the driving system. Moreover, the multi-task, multi-model deployment may increase the computational burden and potentially lead to sub-optimal use of compute.
In contrast to its classical counterpart, an end-to-end autonomous system offers several advantages. (a) The most apparent merit is its simplicity in combining perception, prediction, and planning into a single model that can be jointly trained. (b) The whole system, including its intermediate representations, is optimized towards the ultimate task. (c) Shared backbones increase computational efficiency. (d) Data-driven optimization has the potential to offer emergent abilities that improve the system by simply scaling training resources.
Note that the end-to-end paradigm does not necessarily indicate one black box with only planning/control outputs. It could be modular with intermediate representations and outputs (Fig. 1 (b)) as in the classical approach. In fact, several state-of-the-art systems [1, 2] propose a modular design but optimize all components together to achieve superior performance.
@@ 1.2 Roadmap
Fig. 2 depicts a chronological roadmap of critical achievements in end-to-end autonomous driving, where each part indicates a significant paradigm shift or performance boost. The history of end-to-end autonomous driving dates back to 1988 with ALVINN [3], where the input was two “retinas“ from a camera and a laser range finder, and a simple neural network generated steering output. Bojarski et al. [8] designed a prototype end-to-end CNN system for both simulation and on-road testing, which reestablished this idea in the new era of GPU computing. Notable progress has been achieved with the development of deep neural networks, both in imitation learning [15, 16] and reinforcement learning [4, 17, 18, 19]. The policy distillation paradigm proposed in LBC [5] and related approaches [20, 21, 22, 23] has significantly improved closed-loop performance by mimicking the policy of a well-behaved expert. To enhance generalization ability due to the discrepancy between the expert and learned policy, several papers [10, 24, 25] have proposed aggregating on-policy data [26] during training.
A significant turning point occurred in 2021 for end-to-end autonomous driving. With a diverse range of sensor configurations available within a reasonable computational budget, attention was focused on incorporating more modalities and advanced architectures (e.g., Transformers [27]) to capture global context and representative features, as in TransFuser [6, 28] and many variants [29, 30, 31]. Combined with more insights about the simulation environment, these advanced designs resulted in a substantial performance boost on the closed-loop CARLA benchmark [13]. To improve the interpretability and safety of autonomous systems, approaches such as NEAT [11], NMP [32], and BDD-X [33] explicitly incorporate various auxiliary modules to better supervise the learning process or utilize attention visualization. Recent works prioritize generating safety-critical data [7, 34, 35], pre-training a (large) foundation model or backbone curated for policy learning [36, 12, 37], and advocating a modular end-to-end planning philosophy [2, 38, 39, 1]. Meanwhile, the new and challenging CARLA v2 [13] and nuPlan [14] benchmarks have been introduced to facilitate research into this area.
@@ 1.3 Comparison to Related Surveys
We would like to clarify the difference between our survey and previous related surveys [40, 41, 42, 43, 44, 45, 46, 47, 48]. Some prior surveys [40, 41, 43, 42] cover content similar to ours in the sense of an end-to-end system. However, they do not cover new benchmarks and approaches that arose with the significant recent transition in the field, and place a minor emphasis on frontiers and challenges. The rest of the prior work focuses on specific topics in this domain, such as imitation learning [44, 45, 46] or reinforcement learning [47, 48]. In contrast, our survey provides up-to-date information on the latest developments and techniques in this field, covering a wide span of topics and providing an in-depth discussion of critical challenges.
@@ 2 Methods
This section reviews the fundamental principles behind most existing end-to-end self-driving approaches. Sec. 2.1 discusses methods that use imitation learning and provides details on the two most popular sub-categories, namely behavior cloning and inverse optimal control. Sec. 2.2 summarizes methods that follow the reinforcement learning paradigm.
@@ 2.1 Imitation Learning
Imitation learning (IL), also referred to as learning from demonstrations, trains an agent to learn the optimal policy by imitating the behavior of an expert. IL requires a dataset $D=\{\xi_{i}\}$ containing trajectories collected under the expert’s policy $\pi_{\beta}$, where each trajectory is a sequence of state-action pairs $\{(s_{0},a_{0},s_{1},a_{1},\cdots)\}$. The goal of IL is to learn an agent policy $\pi$ that matches $\pi_{\beta}$. One important and widely used category of IL is behavior cloning (BC) [49], which reduces the problem to supervised learning. Inverse Optimal Control (IOC), also known as Inverse Reinforcement Learning (IRL) [50] is another type of IL method that utilizes expert demonstrations to learn a reward function. We elaborate on these two categories in the following sections.
@@ 2.1.1 Behavior Cloning
In behavior cloning, the objective of matching the agent’s policy with that of the expert is accomplished by minimizing planning loss as a supervised learning problem over the collected dataset: $\mathop{\mathbb{E}}_{(s,a)}\ell(\pi_{\theta}(s),a)$. Here, $\ell(\pi_{\theta}(s),a)$ represents a loss function that measures the distance between the agent action and the expert action.
Early applications of BC for driving tasks [3, 51, 8] utilized an end-to-end neural network to generate control signals from camera inputs. Further enhancements, such as multi-sensor inputs [6, 52], auxiliary tasks [16, 28], and improved expert design [20, 21], have been proposed to enable BC-based end-to-end driving models to handle challenging urban driving scenarios.
Behavior cloning is advantageous due to its simplicity and efficiency, as it does not require hand-crafted reward design, which is crucial for RL. However, there are some common issues associated with behavior cloning. During training, behavior cloning treats each state as independently and identically distributed, resulting in an important problem known as covariate shift. For general IL, several on-policy methods have been proposed to address this issue [53, 54, 55, 26]. In the context of end-to-end autonomous driving, DAgger [26] has been adopted in [10, 5, 56, 25]. Another common problem with behavior cloning is causal confusion, where the imitator exploits and relies on false correlations between certain input components and output signals. This issue has been discussed in the context of end-to-end autonomous driving in [57, 58, 59, 60]. These two challenging problems for imitation learning-based end-to-end autonomous driving are further discussed in Sec. 4.8 and Sec. 4.7, respectively.
@@ 2.1.2 Inverse Optimal Control
Traditional IOC algorithms learn an unknown reward function $R(s,a)$ in a Markov decision process (MDP) from expert demonstrations, where the expert’s reward function can be represented as a linear combination of features [61, 50, 62, 63, 64]. However, in continuous, high-dimensional autonomous driving scenarios, the definition of the reward is implicit and difficult to optimize.
Generative Adversarial Imitation Learning (GAIL) [65, 66, 67] is a specialized approach in IOC that designs the reward function as an adversarial objective to distinguish the expert and learned policies, similar to the concept of generative adversarial networks (GANs) [68]. Recently, several works have proposed optimizing a cost volume or cost function with auxiliary perceptual tasks. Since a cost is an alternative representation of the reward, we classify these methods as belonging to the IOC domain. We define the cost learning framework as follows: end-to-end approaches learn a reasonable cost $c(\cdot)$ in conjunction with other auxiliary tasks, and use simple non-learnable algorithmic trajectory samplers to select the trajectory $\tau^{*}$ with the minimum cost, as illustrated in Fig. 3. Thus, the cost learning paradigm breaks down into two aspects: how to design the cost and how to sample the trajectories to optimize in an end-to-end manner.
Regarding cost design, NMP [32] utilizes a learned cost volume in a bird’s-eye-view (BEV). It also conducts object detection in parallel but does not directly link the cost with the detection outputs. Wang et al. [69] predict future motion for all agents and use the joint energy as an interactive cost to generate the final plan results. [39, 70, 71] propose estimating a set of probabilistic semantic occupancy or freespace layers as an intermediate representation, which provides explicit clues as to where the vehicle should not maneuver to ensure safety. On the other hand, trajectories are typically sampled from a fixed expert trajectory set [1] or processed by parameter sampling with a kinematic model [32, 39, 38, 70]. Then, a max-margin loss is adopted as in classic IOC methods to encourage the expert demonstration to have a minimal cost while others have high costs.
Several challenges still exist with cost learning approaches. In particular, in order to generate more realistic costs, HD maps, auxiliary perception tasks, and multiple sensors are typically incorporated, which increases the difficulty of learning and constructing datasets for multi-modal multi-task frameworks. To tackle this, MP3 [1], ST-P3 [38], and IVMP [72] discard the HD Map inputs used in prior work and utilize the predicted BEV map to calculate the cost of traffic rules, such as staying close to centerlines and avoiding collisions with the road boundary. We provide further discussions in Sec. 4.1 and Sec. 4.4. In general, the aforementioned cost learning approaches significantly enhance the safety and interpretability of decisions made by self-driving cars (see Sec. 4.6), and we believe that the industry-inspired end-to-end system design is a viable approach for real-world applications.
@@ 2.2 Reinforcement Learning
Reinforcement learning (RL) [73] is a field of learning by trial and error. The success of deep Q networks (DQN) [74] in achieving human-level control on the Atari 2600 benchmark [75] has popularized deep reinforcement learning. DQN trains a neural network called the critic (or Q network), which takes as input the current state and an action, and predicts the discounted future reward of that action (when following the same policy afterward). The policy is then implicitly defined by selecting the action with the highest Q value. RL requires an environment that allows potentially unsafe actions to be executed, since it requires exploration (e.g., sometimes executing random actions during data collection). Additionally, RL requires significantly more data to train than supervised learning. For this reason, modern RL approaches often parallelize data collection across multiple environments [76]. Meeting these requirements in real-world cars presents great challenges. Therefore, almost all papers that use RL in autonomous driving have only investigated the technique in simulation. Most works use different extensions of DQN. As of yet, the community has not converged on a specific RL algorithm.
@@
Reinforcement learning has demonstrated successful learning of lane following [4] on a real car on an empty street. Despite this encouraging early result, it must be noted that a similar task was already accomplished by imitation learning three decades prior [3]. To date, no report has shown results for end-to-end training with RL that are competitive with imitation learning. In a direct comparison conducted in conjunction with the release of the CARLA simulator [77], reinforcement learning fell far behind the modular pipeline and end-to-end imitation learning. It is likely that the reason for this failure is that the gradients obtained via RL are insufficient to train the deep perception architectures (ResNet scale) required for driving. The models used in benchmarks like Atari, where RL succeeds, are relatively shallow, consisting of only a few layers [78].
@@
Reinforcement learning has been successfully applied in autonomous driving when combined with supervised learning. Implicit affordances [18] and GRI [19] both pre-train the CNN encoder part of their architecture using supervised learning with auxiliary tasks like semantic segmentation and classification. In the second stage, the pre-trained encoder is frozen, and a shallow policy head is trained on the implicit affordances from the frozen image encoder with a modern version of Q-learning [79]. Both works report state-of-the-art performance on the CARLA leaderboard [13] at the time of writing. Reinforcement learning has also been successfully used to finetune full architectures on CARLA that were pre-trained using imitation learning [17, 80].
@@
RL has also been effectively applied to planning or control tasks where the network has access to privileged simulator information [48, 81, 82]. In the same spirit, RL has been applied to autonomous driving for dataset curation. Roach [21] trains an RL method on privileged BEV semantic segmentation and uses the policy to automatically collect a dataset with which a downstream imitation learning agent is trained. WoR [20] employs a Q function and tabular dynamic programming to generate additional or improved labels for a static dataset.
@@
A future challenge in the field is to transfer the findings from simulation to the real world. In RL, the objective is expressed as a reward function, and most algorithms require these reward functions to be dense and provide feedback at every environment step. Current works typically use simple objectives, such as progress and collision avoidance, and linearly combine them. These simplistic reward functions have been criticized for encouraging risky behavior [81]. Designing or learning better reward functions remains an open problem. Another direction would be to develop RL algorithms that can handle sparse rewards, enabling the optimization of relevant metrics directly. RL can be effectively combined with world models [83, 84, 85], although this presents specific challenges that are discussed in Sec. 4.3. Current RL solutions for autonomous driving rely heavily on low-dimensional representations of the scene, and representation learning is further discussed in Sec. 4.2.

### v2

@@ Abstract
The autonomous driving community has witnessed a rapid growth in approaches that embrace an end-to-end algorithm framework, utilizing raw sensor input to generate vehicle motion plans, instead of concentrating on individual tasks such as detection and motion prediction. End-to-end systems, in comparison to modular pipelines, benefit from joint feature optimization for perception and planning. This field has flourished due to the availability of large-scale datasets, closed-loop evaluation, and the increasing need for autonomous driving algorithms to perform effectively in challenging scenarios. In this survey, we provide a comprehensive analysis of more than 270 papers, covering the motivation, roadmap, methodology, challenges, and future trends in end-to-end autonomous driving. We delve into several critical challenges, including multi-modality, interpretability, causal confusion, robustness, and world models, amongst others. Additionally, we discuss current advancements in foundation models and visual pre-training, as well as how to incorporate these techniques within the end-to-end driving framework. We maintain an active repository that contains up-to-date literature and open-source projects at https://github.com/OpenDriveLab/End-to-end-Autonomous-Driving.
@@ 1 Introduction
Conventional autonomous driving systems adopt a modular design strategy, wherein each functionality, such as perception, prediction, and planning, is individually developed and integrated into onboard vehicles. The planning or control module, responsible for generating steering and acceleration outputs, plays a crucial role in determining the driving experience. The most common approach for planning in modular pipelines involves using sophisticated rule-based designs, which are often ineffective in addressing the vast number of situations that occur on road. Therefore, there is a growing trend to leverage large-scale data and to use learning-based planning as a viable alternative.
@@
In this survey, we conduct an extensive review of this emerging topic. Fig. 1 provides an overview of our work. We begin by discussing the motivation and roadmap for end-to-end autonomous driving systems. End-to-end approaches can be broadly classified into imitation and reinforcement learning, and we give a brief review of these methodologies. We cover datasets and benchmarks for both closed- and open-loop evaluation. We summarize a series of critical challenges, including interpretability, generalization, world models, causal confusion, etc. We conclude by discussing future trends that we think should be embraced by the community to incorporate the latest developments from data engines, and large foundation models, amongst others. Note that this review is mainly orchestrated from a theoretical perspective. Engineering efforts such as version control, unit testing, data servers, data cleaning, software-hardware co-design, etc., play crucial roles in deploying the end-to-end technology. Publicly available information regarding the latest practices on these topics is limited. We invite the community towards more openness in future discussions.
Figure 1: Survey at A Glance. (a) Pipeline and Methods. We define end-to-end autonomous driving as a learning-based algorithm framework with raw sensor input and planning/control output. We deepdive into 270+ papers and categorize into imitation learning (IL) and reinforcement learning (RL). (b) Benchmarking. We group popular benchmarks into closed-loop and open-loop evaluation, respectively. We cover various aspects of closed-loop simulation and the limitations of open-loop evaluation for this problem. (c) Challenges. This is the main section of our work. We list key challenges from a wide range of topics and extensively analyze why these concerns are crucial. Promising resolutions to these challenges are covered as well. (e) Future Trends. We discuss how end-to-end paradigm could benefit by aid of the rapid development of foundation models, visual pre-training, etc. Partial photos by courtesy of online resources.
@@ 1.1 Motivation of an End-to-end System
In the classical pipeline, each model serves a standalone component and corresponds to a specific task (e.g., traffic light detection). Such a design is beneficial in terms of interpretability and ease of debugging. However, since the optimization objectives across modules are different, with detection pursuing mean average precision (mAP) while planning aiming for driving safety and comfort, the entire system may not be aligned with a unified target, i.e., the ultimate planning/control task. Errors from each module, as the sequential procedure proceeds, could be compounded and result in an information loss. Moreover, compared to one end-to-end neural network, the multi-task, multi-model deployment which involves multiple encoders and message transmission systems, may increase the computational burden and potentially lead to sub-optimal use of compute.
In contrast to its classical counterpart, an end-to-end autonomous system offers several advantages. (a) The most apparent merit is its simplicity in combining perception, prediction, and planning into a single model that can be jointly trained. (b) The whole system, including its intermediate representations, is optimized towards the ultimate task. (c) Shared backbones increase computational efficiency. (d) Data-driven optimization has the potential to improve the system by simply scaling training resources.
Note that the end-to-end paradigm does not necessarily indicate one black box with only planning/control outputs. It could have intermediate representations and outputs (Fig. 1 (b)) as in classical approaches. In fact, several state-of-the-art systems [1, 2] propose a modular design but optimize all components together to achieve superior performance.
@@ 1.2 Roadmap
Fig. 2 depicts a chronological roadmap of critical achievements in end-to-end autonomous driving, where each part indicates an essential paradigm shift or performance boost. The history of end-to-end autonomous driving dates back to 1988 with ALVINN [3], where the input was two “retinas“ from a camera and a laser range finder, and a simple neural network generated steering output. NVIDIA designed a prototype end-to-end CNN system, which reestablished this idea in the new era of GPU computing [8]. Notable progress has been achieved with the development of deep neural networks, both in imitation learning [15, 16] and reinforcement learning [4, 17, 18, 19]. The policy distillation paradigm proposed in LBC [5] and related approaches [20, 21, 22, 23] has significantly improved closed-loop performance by mimicking a well-behaved expert. To enhance generalization ability due to the discrepancy between the expert and learned policy, several papers [10, 24, 25] have proposed aggregating on-policy data [26] during training.
A significant turning point occurred around 2021. With diverse sensor configurations available within a reasonable computational budget, attention was focused on incorporating more modalities and advanced architectures (e.g., Transformers [27]) to capture global context and representative features, as in TransFuser [6, 28] and many variants [29, 30, 31]. Combined with more insights about the simulation environment, these advanced designs resulted in a substantial performance boost on the CARLA benchmark [13]. To improve the interpretability and safety of autonomous systems, approaches [11, 32, 33] explicitly involve various auxiliary modules to better supervise the learning process or utilize attention visualization. Recent works prioritize generating safety-critical data [7, 34, 35], pre-training a foundation model or backbone curated for policy learning [36, 12, 37], and advocating a modular end-to-end planning philosophy [2, 38, 39, 1]. Meanwhile, the new and challenging CARLA v2 [13] and nuPlan [14] benchmarks have been introduced to facilitate research into this area.
@@ 1.3 Comparison to Related Surveys
We would like to clarify the difference between our survey and previous related surveys [40, 41, 42, 43, 44, 45, 46, 47, 48]. Some prior surveys [40, 41, 43, 42] cover content similar to ours in the sense of an end-to-end system. However, they do not cover new benchmarks and approaches that arose with the significant recent transition in the field, and place a minor emphasis on frontiers and challenges. The others focus on specific topics in this domain, such as imitation learning [44, 45, 46] or reinforcement learning [47, 48]. In contrast, our survey provides up-to-date information on the latest developments in this field, covering a wide span of topics and providing in-depth discussions of critical challenges.
@@ 2 Methods
This section reviews fundamental principles behind most existing end-to-end self-driving approaches. Sec. 2.1 discusses methods using imitation learning and provides details on the two most popular sub-categories, namely behavior cloning and inverse optimal control. Sec. 2.2 summarizes methods that follow the reinforcement learning paradigm.
@@ 2.1 Imitation Learning
Imitation learning (IL), also referred to as learning from demonstrations, trains an agent to learn the policy by imitating the behavior of an expert. IL requires a dataset $D=\{\xi_{i}\}$ containing trajectories collected under the expert’s policy $\pi_{\beta}$, where each trajectory is a sequence of state-action pairs. The goal of IL is to learn an agent policy $\pi$ that matches $\pi_{\beta}$.
The policy $\pi$ can output planned trajectories or control signals. Early works usually adopt control outputs, due to the ease of collection. However, predicting controls at different steps could lead to discontinuous maneuvers and the network inherently specializes to the vehicle dynamics which hinders generalization to other vehicles. Another genre of works predicts waypoints. It considers a relatively longer time horizon. Meanwhile, converting trajectories for vehicles to track into control signals needs additional controllers, which is non-trivial and involves vehicle models and control algorithms. Since no clear performance gap has been observed between these two paradigms, we do not differentiate them explicitly in this survey. An interesting and more in-depth discussion can be found in [22].
One widely used category of IL is behavior cloning (BC) [49], which reduces the problem to supervised learning. Inverse Optimal Control (IOC), also known as Inverse Reinforcement Learning (IRL) [50] is another type of IL method that utilizes expert demonstrations to learn a reward function. We elaborate on these two categories below.
@@ 2.1.1 Behavior Cloning
In BC, matching the agent’s policy with the expert’s is accomplished by minimizing planning loss as supervised learning over the collected dataset: $\mathop{\mathbb{E}}_{(s,a)}\ell(\pi_{\theta}(s),a)$. Here, $\ell(\pi_{\theta}(s),a)$ represents a loss function that measures the distance between the agent action and the expert action.
Early applications of BC for driving [3, 51, 8] utilized an end-to-end neural network to generate control signals from camera inputs. Further enhancements, such as multi-sensor inputs [6, 52], auxiliary tasks [16, 28], and improved expert design [21], have been proposed to enable BC-based end-to-end driving models to handle challenging urban scenarios.
BC is advantageous due to its simplicity and efficiency, as it does not require hand-crafted reward design, which is crucial for RL. However, there are some common issues. During training, it treats each state as independently and identically distributed, resulting in an important problem known as covariate shift. For general IL, several on-policy methods have been proposed to address this issue [53, 54, 55, 26]. In the context of end-to-end autonomous driving, DAgger [26] has been adopted in [10, 5, 56, 25]. Another common problem with BC is causal confusion, where the imitator exploits and relies on false correlations between certain input components and output signals. This issue has been discussed in the context of end-to-end autonomous driving in [57, 58, 59, 60]. These two challenging problems are further discussed in Sec. 4.9 and Sec. 4.8, respectively.
@@ 2.1.2 Inverse Optimal Control
Traditional IOC algorithms learn an unknown reward function $R(s,a)$ from expert demonstrations, where the expert’s reward function can be represented as a linear combination of features [61, 50, 62, 63, 64]. However, in continuous, high-dimensional autonomous driving scenarios, the definition of the reward is implicit and difficult to optimize.
Generative adversarial imitation learning [65, 66, 67] is a specialized approach in IOC that designs the reward function as an adversarial objective to distinguish the expert and learned policies, similar to the concept of generative adversarial networks [68]. Recently, several works propose optimizing a cost volume or cost function with auxiliary perceptual tasks. Since a cost is an alternative representation of the reward, we classify these methods as belonging to the IOC domain. We define the cost learning framework as follows: end-to-end approaches learn a reasonable cost $c(\cdot)$ and use algorithmic trajectory samplers to select the trajectory $\tau^{*}$ with the minimum cost, as illustrated in Fig. 3.
Regarding cost design, it has representations including a learned cost volume in a bird’s-eye-view (BEV) [32], joint energy calculated from other agents’ future motion [69], or a set of probabilistic semantic occupancy or freespace layers [39, 70, 71]. On the other hand, trajectories are typically sampled from a fixed expert trajectory set [1] or processed by parameter sampling with a kinematic model [32, 39, 38, 70]. Then, a max-margin loss is adopted as in classic IOC methods to encourage the expert demonstration to have a minimal cost while others have high costs.
Several challenges exist with cost learning approaches. In particular, in order to generate more realistic costs, HD maps, auxiliary perception tasks, and multiple sensors are typically incorporated, which increases the difficulty of learning and constructing datasets for multi-modal multi-task frameworks. Nevertheless, the aforementioned cost learning methods significantly enhance the safety and interpretability of decisions (see Sec. 4.6), and we believe that the industry-inspired end-to-end system design is a viable approach for real-world applications.
@@ 2.2 Reinforcement Learning
Reinforcement learning (RL) [72, 73] is a field of learning by trial and error. The success of deep Q networks (DQN) [74] in achieving human-level control on the Atari benchmark [75] has popularized deep RL. DQN trains a neural network called the critic (or Q network), which takes as input the current state and an action, and predicts the discounted return of that action. The policy is then implicitly defined by selecting the action with the highest predicted return.
@@
RL requires an environment that allows potentially unsafe actions to be executed, to collect novel data (e.g., via random actions). Additionally, RL requires significantly more data to train than IL. For this reason, modern RL methods often parallelize data collection across multiple environments [76]. Meeting these requirements in the real world presents great challenges. Therefore, almost all papers that use RL in driving have only investigated the technique in simulation. Most use different extensions of DQN. The community has not yet converged on a specific RL algorithm.
@@
RL has successfully learned lane following on a real car on an empty street [4]. Despite this encouraging result, it must be noted that a similar task was already accomplished by IL three decades prior [3]. To date, no report has shown results for end-to-end training with RL that are competitive with IL. The reason for this failure likely is that the gradients obtained via RL are insufficient to train deep perception architectures (i.e., ResNet) required for driving. Models used in benchmarks like Atari, where RL succeeds, are relatively shallow, consisting of only a few layers [77].
@@
RL has been successfully applied in end-to-end driving when combined with supervised learning (SL). Implicit affordances [18, 19] pre-train the CNN encoder using SL with tasks like semantic segmentation. In the second stage, this encoder is frozen, and a shallow policy head is trained on the features from the frozen encoder with a modern version of Q-learning [78]. RL can also be used to finetune full networks that were pre-trained using IL [17, 79].
@@
RL can also been effectively applied, if the network has access to privileged simulator information. [48, 80, 81]. Privileged RL agents can be used for dataset curation. Roach [21] trains an RL agent on privileged BEV semantic maps and uses the policy to automatically collect a dataset with which a downstream IL agent is trained. WoR [20] employs a Q-function and tabular dynamic programming to generate additional or improved labels for a static dataset.
@@
A challenge in the field is to transfer the findings from simulation to the real world. In RL, the objective is expressed as reward functions, and many algorithms require them to be dense and provide feedback at each environment step. Current works typically use simple objectives, such as progress and collision avoidance. These simplistic designs potentially encourage risky behaviors [80]. Devising or learning better reward functions remains an open problem. Another direction would be to develop RL algorithms that can handle sparse rewards, enabling the optimization of relevant metrics directly. RL can be effectively combined with world models [82, 83, 84], though this presents specific challenges (See Sec. 4.3). Current RL solutions for driving rely heavily on low-dimensional representations of the scene, and this issue is further discussed in Sec. 4.2.2.